A hydrogen atom is 5.3 × 10−10 m in radius. Use the uncertainty principle to estimate the minimum energy an electron can have in this atom. 


1
Expert's answer
2022-03-22T15:48:52-0400

Answer

According to uncertainty

"P=\frac{\hbar}{r}\\=2.11*10^{-20}kg-m/s"

So minimum energy

"E=\sqrt{p^2c^2+m^2c^4}"

Putting all values

E="3.3*10^{-12}J"